# Revisiting Automatic Data Curation for Vision Foundation Models in Digital Pathology

### v1

@@ 1 Introduction
Data curation plays a crucial role in training FMs, leading to improved performance [17]. Strategies such as data pruning [20, 18], active learning [10, 22], and nearest neighbor search [17] have proven effective. Recently, clustering-based methods have emerged as competitive alternatives to traditional supervised algorithms, while being fully automated [21]. In contrast, data curation in DP largely relies on expert annotations that provide high-level information at the WSI level (e.g., cancer and tissue types [14]), while potentially overlooking tile-level granularity. To this end, a semi-automatic strategy has been proposed recently, combining WSI labels and weak tile-level supervision based on color statistics [8]. Despite this progress, the need for expert intervention still limits scalability, underscoring the necessity of a fully automated data curation framework tailored to the heterogeneity and annotation challenges of DP.
@@
To summarize, our contributions are three-fold: 1) we present, to the best of our knowledge, the first fully automated data curation scheme for FM training in DP; 2) we identify batch stratification, alongside hierarchical clustering-based curation, as the key to data distribution uniformization and improved downstream performance; 3) we demonstrate the effectiveness of our method through improved performance on both region of interest (RoI)- and WSI-level benchmarks. Our code can be found on GitHub [^1] https://github.com/swiss-ai/health-pathology/tree/main.
@@ 2.1 Hierarchical Clustering for Automatic Data Curation
Inspired by [21], we first extract tile embeddings from the pre-training dataset using a well-established FM in DP (i.e., UNI [6]) and construct a hierarchical clustering tree in a bottom-up manner. At the bottom level, K-means is applied to all tile embeddings, generating more clusters and of relatively smaller volumes in dense areas than in sparse ones. Data imbalance is artificially reduced, while the average nearest-neighbor distance is increased, by selecting the cluster centroids as the new data distribution. We recursively apply K-means to these centroids with an exponentially decreasing number of clusters, ensuring more uniform cluster volumes at each hierarchical level. This recursion provably yields clusters that distribute more uniformly over the data support as one ascends the hierarchy [17]. Next, a top-down sampling strategy is applied to sample curated subsets of size $N$. First, sample sizes are allocated to the $k$ top-level clusters $\{c_{i}\}_{i=1}^{k}$ of respective total sizes $\{s_{i}\}^{k}_{i=1}$, by finding the size $n$ which minimizes $|N-\sum_{i}^{k}min(n,s_{i})|$ through a binary search in $\{0,\ldots,N\}$. This process is then repeated for each cluster $c_{i}$ independently, distributing $min(n,s_{i})$ samples across its respective downstream clusters. These two steps define a recursion applied until reaching the bottom-level clusters, where the allocated points are sampled randomly. Overall, this method allows us to efficiently sample subsets of great diversity, while controlling the trade-off between balance and size.
@@
We note that T2 closely mirrors the hyperparameters in [21], with a higher top-level cluster count (equals to the global batch size, see Section 2.2) and branch expansion rates, whereas T1 has fewer top-level clusters, following insights from expert annotations [14], thereby requiring greater compression at each level. In Figure 2, we use these trees to analyze the data distribution and illustrate the effects of hierarchical sampling. Figure 2(a) presents UMAP embeddings of top-level centroids from T1, with hyperparameters chosen to best preserve local relative positions. Centroids are colored based on their effective size without sampling. The results indicate that top-level clusters indeed cover rather uniformly our distribution and that it has heavy tails with 2 modes encompassing approximately two-thirds of the dataset. The results from sampling 10% of the data (Figure 2(b)) show that while global diversity is preserved, the curated subset remains slightly imbalanced. We further analyze the trade-offs between balance and size via the total variation distance (TV) between sampled cluster proportions ($p$) and uniform weights ($p_{unif}$) across two clustering levels. Results shown in Figure 2(c) and (d) reveal that the large imbalance in our dataset implies a significant imbalance for all curated subsets with more than 1% data, since the smallest top-level clusters are depleted during sampling. We also observe that sampling from any level results in a nearly maximal imbalance at the other level studied, as densest level 4 clusters contain more clusters at level 3.
@@ 2.2 Effective Self-supervised Learning on Curated Data
After the data curation, we proceed to the self-supervised training of FMs using the curated datasets. We follow existing works on FMs in DP [23, 24, 6, 8] and adopt DINOv2 [17] as our SSL framework. We utilize large Vision Transformers (ViT-L) [9] and train it for 170 thousand iterations with a global batch size of 2048, corresponding to one complete pass over the full dataset (350 million tiles). Typically, these batches are randomly sampled from the pre-training dataset (curated or not), as in most FMs training with automatic data curation [21] or without it [17, 23, 24, 6]. Instead, we propose to structure these batches during training by stratifying them based on the tree clusters from which the tiles were sampled. In particular, each batch is constructed to contain an equal number of tiles from each top-level cluster, leveraging the inherent balancing of curation clusters to guarantee uniform coverage of the data distribution within batches. This approach ensures that differently populated clusters are equally represented during training, prompting models to continually adapt to diverse samples while mitigating biases introduced by dataset imbalance. To further promote intra-cluster diversity in batch construction, we track the samples encountered during training and prioritize those that have been less frequently sampled: for each batch and cluster, we sample exclusively from the least observed tiles within that batch and cluster. This procedure guarantees that every tile in the curated dataset is observed multiple times during training.
@@ 3 Experiments
Interpretability of hierarchical clusters. To study the interpretability of hierarchical clusters, we analyze the 1% curated subset using T1 at level 4. We first conduct visual inspections of tile images in the clusters, an example of which is shown in the supplementary video. We then follow [25] to quantitatively report our findings, by extracting basic statistics for handcrafted single-cell features (color intensities, textures, morphologies, spatial arrangements) and including class densities from CellViT [13] and extracellular matrix descriptors. We observe that $\sim$50highest-level clusters mainly contain tiles of healthy tissues from GTEx and a few TCGA tiles without neoplastic cells, while the two largest clusters are mainly cancerous TCGA tiles with $\sim$50cells neoplastic or inflammatory and the remainder epithelial or connective tissue cells. Next, we cluster the tiles using handcrafted features by applying UMAP, followed by K-means with the same number of clusters as T1’s top level. The resulting clusters are then compared to those obtained from hierarchical clustering on raw embeddings using the Adjusted Rand Index (ARI). Our experiments show that both types of clusters match weakly with ARI scores of 7.7% at level 4 and 5.5% at level 3. Moreover, restricting the same process to specific types of handcrafted features reveals that cell colors and spatial arrangements are relatively the most discriminative factors across hierarchical clusters. Overall, these results suggest that while handcrafted features explain the hierarchical clusters to some extent, they do not fully capture the underlying patterns dictating this organization.
@@
Evaluation framework. In the following, we evaluate different data curation settings at both RoI- and WSI-level. Specifically, 8 RoI-level tasks from independent cohorts are considered, including lung adenocarcinoma classification (LUAD [11]), colorectal tissue and polyp classification (CRC [16], UniToPatho [2], Chaoyang [26]), breast cancer subtyping (BRACS [3]) and tissue classification (BACH [1], BreakHis [19]), and lymph node metastatis classification (PCAM). All RoI images are resized to 224×224 pixels before embedding extraction. We perform linear probing over 1000 nonparametric bootstrap iterations on embeddings extracted from each encoder. At the WSI-level, we assess 9 clinically relevant tasks [7], including breast cancer (BCa) detection, and biomarker prediction for Estrogen Receptor (ER), Progesterone Receptor (PR) and Human Epidermal Growth Factor Receptor (HER2) and breast Homologous Repair Deficiency status (HRD). Further tasks include prediction of lung Epidermal Growth Factor Receptor (EGFR) mutation status and immunotherapy (IO) response in lung cancer patients, and, lastly, detection of inflammatory bowel disease (IBD) versus normal mucosa samples. We represent each WSI as a set of tile embeddings composing its segmented tissue regions and train ABMIL [15] models over 20 Monte Carlo cross-validation runs [15]. We report average balance accuracy (bACC, in %) and area under the receiver operating characteristic curve (AUC, in %) for RoI- and WSI-level tasks, respectively [6, 7].
@@
Main results. The performance on the RoI- and WSI-level benchmarks are presented in Tables 1 and 2, respectively. We can observe that our approach, T1-BS, achieves the highest average performance across both benchmarks, followed by S-BS and F-BR. In contrast, T1-BR and S-BR yield the lowest average performance. These results emphasize the effectiveness of jointly applying automatic data curation at the tile level and stratified batch sampling strategies. Nevertheless, rankings per task still vary significantly. For further analysis, we note that our benchmarks predominantly consist of images related to breast, colon, and lung tissues. Thus, we can expect any inherent bias towards these tissue types in benchmarked methods could lead to improved performance. To better disentangle model performance from such biases, we first provide a detailed breakdown of the dataset compositions used during training. In the full dataset, breast, colon, and lung tissues account for 5.2%, 5.1%, and 6.7% of tiles, with 61% coming from TCGA, which are more likely cancer tissues. In the supervised curation subset, these proportions shift to 6.1%, 3.2%, and 9.3%, with 80% from TCGA, dispersed across 58 of 266 classes. The hierarchical clustering subset contains 2.7%, 3.9%, and 3.9%, with 53% from TCGA, covering all level-4 clusters. From a global perspective, both T1-BS and T1-BR utilize on average the smallest proportions of tiles from these tissues while maintaining competitive performance, underscoring the effectiveness of automatic data curation in enhancing diversity in sampling. Among the most comparable settings, e.g. T1-BS and S-BS, we observe strong correlations between performance and both tile proportions and their origins. Both methods use similar tile proportions from GTex for breast and lung tiles, but S-BS uses increasingly more TGCA tiles for each tissue. However, T1-BS outperforms S-BS on RoI-level tasks related to breast tissue, achieving an average balanced accuracy of 87.2% compared to 84.7% for S-BS, while demonstrating slightly better performance on the LUAD lung dataset. Notably, T1-BS leads to the largest improvement on the most challenging task, BRACS, with 7 cancer subtypes. For colon tissues, S-BS uses fewer tiles than T1-BS yet their average performance remains comparable at 71.1% and 70.9%, respectively. T1-BS actually tends to outperform S-BS in benchmarks with greater tissue diversity, achieving superior performance on CRC (9 classes) while performing similarly or slightly worse on UniToPatho (6 classes) and Chaoyang (4 classes), suggesting its potential for training a better generalist FM.
@@
Site 1 Site 2 Site 1
@@ 4.0.1 Acknowledgements
We gratefully acknowledge funding from the ETH AI Center, the Idiap Research Institute, and the Swiss Federal Institutes of Technology strategic focus on personalized health and related technologies, as well as the EPFL and ETH core funding. We also acknowledge funding from the Novartis Foundation for Medical-Biological Research (grant no. 22B104) and from the Swiss AI Initiative through a grant by the Swiss National Supercomputing Centre (CSCS) under project ID a02 on Alps.

### v2

@@ 1 Introduction
Data curation plays a crucial role in training FMs, leading to improved performance [17]. Strategies such as data pruning [20, 18], active learning [10, 22], and nearest neighbor search [17] have proven effective. Recently, clustering-based methods have emerged as competitive alternatives to traditional supervised algorithms, while being fully automated [21]. In contrast, data curation in DP largely relies on expert annotations that provide high-level information at the WSI level (e.g.,, cancer and tissue types [14]), while potentially overlooking tile-level granularity. To this end, a semi-automatic strategy has been proposed recently, combining WSI labels and weak tile-level supervision based on color statistics [8]. Despite this progress, the need for expert intervention still limits scalability, underscoring the necessity of a fully automated data curation framework tailored to the heterogeneity and annotation challenges of DP.
@@
To summarize, our contributions are three-fold: 1) we present, to the best of our knowledge, the first fully automated data curation scheme for FM training in DP; 2) we identify batch stratification, alongside hierarchical clustering-based curation, as the key to data distribution uniformization and improved downstream performance; 3) we demonstrate the effectiveness of our method through improved performance on both region of interest (RoI)- and WSI-level benchmarks. Our code and curated data can be found on GitHub [^1] https://github.com/swiss-ai/patho-ssl-data-curation.
@@ 2.1 Hierarchical Clustering for Automatic Data Curation
These two steps define a recursion applied until reaching the bottom-level clusters, where the allocated points are sampled randomly. Inspired by [21], we first extract tile embeddings from the pre-training dataset using a well-established FM in DP (i.e., UNI [6]) and construct a hierarchical clustering tree in a bottom-up manner. At the bottom level, K-means is applied to all tile embeddings, generating more clusters and of relatively smaller volumes in dense areas than in sparse ones. Data imbalance is artificially reduced, while the average nearest-neighbor distance is increased, by selecting the cluster centroids as the new data distribution. We recursively apply K-means to these centroids with an exponentially decreasing number of clusters, ensuring more uniform cluster volumes at each hierarchical level. This recursion provably yields clusters that distribute more uniformly over the data support as one ascends the hierarchy [17]. Remark that both K-means++ and the resampling strategies proposed in [21] are utilized in practice to improve the final hierarchical clustering. Next, a top-down sampling strategy is applied to sample curated subsets of size $N$. First, sample sizes are allocated to the $k$ top-level clusters $\{c_{i}\}_{i=1}^{k}$ of respective total sizes $\{s_{i}\}^{k}_{i=1}$, by finding the size $n$ which minimizes $|N-\sum_{i}^{k}min(n,s_{i})|$ through a binary search in $\{0,\ldots,N\}$. This process is then repeated for each cluster $c_{i}$ independently, distributing $min(n,s_{i})$ samples across its respective downstream clusters. Overall, this method allows us to efficiently sample subsets of great diversity, while controlling the trade-off between balance and size.
@@
We note that T2 closely mirrors the hyperparameters in [21], with a higher top-level cluster count (equals to the global batch size, see Section 2.2) and branch expansion rates, whereas T1 has fewer top-level clusters, following insights from expert annotations [14], thereby requiring greater compression at each level. In Figure 2, we use these trees to analyze the data distribution and illustrate the effects of hierarchical sampling. Figure 2(a) presents UMAP embeddings of top-level centroids from T1, with hyperparameters chosen to best preserve local relative positions. Centroids are colored based on their effective size without sampling. The results indicate that top-level clusters indeed cover rather uniformly our distribution and that it has heavy tails with 2 modes encompassing approximately two-thirds of the dataset. The results from sampling 10% of the data (Figure 2(b)) show that while global diversity is preserved, the curated subset remains slightly imbalanced. We further analyze the trade-offs between balance and size via the total variation distance (TV) between sampled cluster proportions ($p$) and uniform weights ($p_{unif}$) across two clustering levels. Results shown in Figure 2(c) and (d) reveal that the large imbalance in our dataset implies a significant imbalance for all curated subsets with more than 1% data, since the smallest top-level clusters are depleted during sampling. We also observe that sampling from any level results in a nearly maximal imbalance at the other level studied, as densest level 4 clusters contain more clusters at level 3. This implies, for instance, that uniform sampling of a small amount of data at level 4, e.g., with TV near 0 at this level, propagates more samples in fewer level 3 clusters within sparser level 4 clusters, and vice versa in denser level 4 clusters, resulting in a high TV at level 3.
@@ 2.2 Effective Self-supervised Learning on Curated Data
After the data curation, we proceed to the self-supervised training of FMs using the curated datasets. We follow existing works on FMs in DP [23, 24, 6, 8] and adopt DINOv2 [17] as our SSL framework. We utilize large Vision Transformers (ViT-L) [9], trained on the curated 10% subset for 170 thousand iterations with a global batch size of 2048, equivalent to one complete pass over the full dataset (350 million tiles). Typically, these batches are randomly sampled from the pre-training dataset (curated or not), as in most FMs training with automatic data curation [21] or without it [17, 23, 24, 6]. Instead, we propose to structure these batches during training by stratifying them based on the tree clusters from which the tiles were sampled. In particular, each batch is constructed to contain an equal number of tiles from each top-level cluster, leveraging the inherent balancing of curation clusters to guarantee uniform coverage of the data distribution within batches. This approach ensures that differently populated clusters are equally represented during training, prompting models to continually adapt to diverse samples while mitigating biases introduced by dataset imbalance. To further promote intra-cluster diversity in batch construction, we track the samples encountered during training and prioritize those that have been less frequently sampled: for each batch and cluster, we sample exclusively from the least observed tiles within that batch and cluster. This procedure guarantees that every tile in the curated dataset is observed multiple times during training.
@@ 3 Experiments
Interpretability of hierarchical clusters. To study the interpretability of hierarchical clusters, we analyze the 1% curated subset using T1 at level 4. We first conduct visual inspections of tile images in the clusters, an example of which is shown in the supplementary video. We then follow [25] to quantitatively report our findings, by extracting basic statistics for handcrafted single-cell features (color intensities, textures, morphologies, spatial arrangements) and including class densities from CellViT [13] and extracellular matrix descriptors. We observe that $\sim$50highest-level clusters mainly contain tiles of healthy tissues from GTEx and a few TCGA tiles without neoplastic cells, while the two largest clusters are mainly cancerous TCGA tiles with $\sim$50cells neoplastic or inflammatory and the remainder epithelial or connective tissue cells.
Next, we cluster the tiles using handcrafted features by applying UMAP, followed by K-means with the same number of clusters as T1’s top level. The resulting clusters are then compared to those obtained from hierarchical clustering on raw embeddings using the Adjusted Rand Index (ARI). Our experiments show that both types of clusters match weakly with ARI scores of 7.7% at level 4 and 5.5% at level 3. Moreover, restricting the same process to specific types of handcrafted features reveals that cell colors and spatial arrangements are relatively the most discriminative factors across hierarchical clusters. Overall, these results suggest that while handcrafted features explain the hierarchical clusters to some extent, they do not fully capture the underlying patterns dictating this organization.
@@
Evaluation framework. In the following, we evaluate different data curation settings at both RoI- and WSI-level. Specifically, 8 RoI-level tasks from independent cohorts are considered, including lung adenocarcinoma classification (LUAD [11]), colorectal tissue and polyp classification (CRC [16], UniToPatho [2], Chaoyang [26]), breast cancer subtyping (BRACS [3]) and tissue classification (BACH [1], BreakHis [19]), and lymph node metastatis classification (PCAM). All RoI images are resized to 224×224 pixels before embedding extraction. We perform linear probing over 1000 nonparametric bootstrap iterations on embeddings extracted from each encoder. At the WSI-level, we assess 9 clinically relevant tasks [7] from two different clinical centers (MSHS and MSKCC), including breast cancer (BCa) detection, and biomarker prediction for Estrogen Receptor (ER), Progesterone Receptor (PR) and Human Epidermal Growth Factor Receptor (HER2) and breast Homologous Repair Deficiency status (HRD). Further tasks include prediction of lung Epidermal Growth Factor Receptor (EGFR) mutation status and immunotherapy (IO) response in lung cancer patients, and, lastly, detection of inflammatory bowel disease (IBD) versus normal mucosa samples. We represent each WSI as a set of tile embeddings composing its segmented tissue regions and train attention-based multiple instance learning [15] models over 20 Monte Carlo cross-validation runs [15]. We report average balance accuracy (bACC, in %) and area under the receiver operating characteristic curve (AUC, in %) for RoI- and WSI-level tasks, respectively [6, 7].
@@
Main results. The performance on the RoI- and WSI-level benchmarks are presented in Tables 1 and 2, respectively. We can observe that our approach, T1-BS, achieves the highest average performance across both benchmarks, followed by S-BS and F-BR. In contrast, T1-BR and S-BR yield the lowest average performance. These results emphasize the effectiveness of jointly applying automatic data curation at the tile level and stratified batch sampling strategies. Nevertheless, rankings per task still vary significantly. For further analysis, we note that our benchmarks predominantly consist of images related to breast, colon, and lung tissues. Thus, we can expect any inherent bias towards these tissue types in benchmarked methods could lead to improved performance. To better disentangle model performance from such biases, we first provide a detailed breakdown of the dataset compositions used during training. In the full dataset, breast, colon, and lung tissues account for 5.2%, 5.1%, and 6.7% of tiles, with 61% coming from TCGA, which are more likely cancer tissues. In the supervised curation subset, these proportions shift to 6.1%, 3.2%, and 9.3%, with 80% from TCGA, dispersed across 58 of 266 classes. The hierarchical clustering subset contains 2.7%, 3.9%, and 3.9%, with 53% from TCGA, covering all level-4 clusters. From a global perspective, both T1-BS and T1-BR utilize on average the smallest proportions of tiles from these tissues while maintaining competitive performance, underscoring the effectiveness of automatic data curation in enhancing diversity in sampling. Among the most comparable settings, e.g., T1-BS and S-BS, we observe strong correlations between performance and both tile proportions and their origins. Both methods use similar tile proportions from GTex for breast and lung tiles, but S-BS uses increasingly more TGCA tiles for each tissue. However, T1-BS outperforms S-BS on RoI-level tasks related to breast tissue, achieving an average balanced accuracy of 87.2% compared to 84.7% for S-BS, while demonstrating slightly better performance on the LUAD lung dataset. Notably, T1-BS leads to the largest improvement on the most challenging task, BRACS, with 7 cancer subtypes. For colon tissues, S-BS uses fewer tiles than T1-BS yet their average performance remains comparable at 71.1% and 70.9%, respectively. T1-BS actually tends to outperform S-BS in benchmarks with greater tissue diversity, achieving superior performance on CRC (9 classes) while performing similarly or slightly worse on UniToPatho (6 classes) and Chaoyang (4 classes), suggesting its potential for training a better generalist FM.
@@
MSHS MSKCC MSHS
@@ 4.0.1 Acknowledgements
We gratefully acknowledge funding from the ETH AI Center, the Idiap Research Institute, and the Swiss Federal Institutes of Technology strategic focus on personalized health and related technologies, as well as the EPFL and ETH core funding. We acknowledge funding from the Novartis Foundation for Medical-Biological Research (grant no. 22B104) and from the Swiss AI Initiative through a grant by the Swiss National Supercomputing Centre (CSCS) under project ID a02 on Alps.